# Multimodal Outer Arithmetic Block Dual Fusion of Whole Slide Images and Omics Data for Precision Oncology

### v1

@@ 2.2.1 Omic Encoder
To construct the omic encoder, we followed the established practice of using a Self-Normalizing Neural (SNN) [32] consisting of two fully connected layers, where each layer applies an Exponential Linear Unit (ELU) activation function followed by Alpha Dropout (0.25) that compress the 8K CpG sites into an encoded omic feature ${\mathbf{o}_{i}}\in\mathbb{R}^{d^{o}}$ ${d^{o}}=256$. Note that for comparison studies [22, 13, 20, 18], we used the same encoder tokenizing the genes into a group of pathways to draw a fair comparison.
@@ 2.2.2 Whole Slide Image Encoder
For a WSI $\mathbf{X}$, we extract a collection of patches, represented as ${\mathbf{x}_{i0},\mathbf{x}_{i1},\ldots,\mathbf{x}_{ij}}$; where $i$ denotes the patch index and $j$ indicates the variable number of patches across different slides. We extract non-overlapping patches from tissue areas at a 20× magnification (about 0.5 µm/pixel resolution). Subsequently, we utilized a SOTA vision-only encoder [11] UNI to obtain the patch embeddings ${\mathbf{e}_{ij}}$. Using UNI ${f_{enc}(.)}$, we derive a set of low-dimensional patch embeddings for each patient, where ${\mathbf{e}_{ij}}={f_{enc}}(\mathbf{x}_{ij})\in\mathbb{R}^{d^{e}}$, ${d^{e}}=1024$, serving as input to our pipeline.
@@ 2.2.3 Fusion Stages
Early Fusion, given a matrix of patch embeddings ${\mathbf{e}_{ij}}$ and the DNA feature vector ${\mathbf{o}_{i}}$ for patient $i$, the encoded omic feature vector is cloned to match the number of patches in the WSI. This results in a tensor of shape $(N_{i},d_{o})$, where $N_{i}$ is the number of patches in WSI $i$, and $d_{o}$ is the dimension of the DNA feature. The combined feature set for patient $i$ is represented as:
@@
This operation is performed for all patches $j$ of the $i$th patient. By incorporating omic features $\mathbf{o}_{i}$ into patch embeddings via early fusion, we enrich the representation with complementary molecular information. Leveraging an Attention-based Deep Multiple Instance Learning (ABMIL) approach[33], we capture patch-level discriminative features that synergistically combine molecular and morphological insights, enabling more precise identification of critical regions. The resulting embedding is projected to a slide-level representation $\mathbf{W}_{i}\in\mathbb{R}^{256}$, dimensionally aligned with the original omic feature $\mathbf{o}_{i}$ to facilitate subsequent late fusion.
@@
Late Fusion, motivated by [13] in modeling the interaction between omic to histology, histology to omic, and omic to omic, we mimic a similar behavior by employing, for the first time in journal form, a novel multimodal outer arithmetic fusion block (MOAB) [28] within a dual fusion approach, marking a new direction in combining MOAB with dual fusion to enhance modality integration. MOAB inputs ($\mathbf{W}_{i}$,$\mathbf{o}_{i}$) will be fused through four operations: outer product, outer division, outer subtraction, and outer addition. MOAB extracts various interactions while simultaneously preserving the $\mathbf{W}_{i}$ input feature by appending one to each input embedding when performing the outer product and division fusion and zero in the case of the outer subtraction and addition fusion. To simplify this, given the two embeddings $\mathbf{W}_{i}$,$\mathbf{o}_{i}$ $\mathbf{\mathbf{W}_{i}}\in\mathbb{R}^{N\times 1}$ and $\mathbf{\mathbf{o}_{i}}\in\mathbb{R}^{M\times 1}$. Initially, we append a 1 to each embedding, i.e., $\mathbf{W}_{i1}=[1;\mathbf{W}_{i}]$ and $\mathbf{o}_{j1}=[1;\mathbf{o}_{j}]$. Their outer product is defined as
$(\mathbf{\mathbf{W}}_{i1}\otimes\mathbf{o}_{j1})_{ij}=\mathbf{W}_{1i}*\mathbf{o}_{1j},$ (3)
@@
for $i\in[1...N+1]$ and $j\in[1...M+1]$ yielding a $(N+1)\times(M+1)$ matrix that intermingles every element of $\mathbf{W}_{i1}$ with every element of $\mathbf{o}_{j1}$. The appended 1 in both $\mathbf{W}_{i1}$ and $\mathbf{o}_{j1}$ ensures the original unimodal features $\mathbf{\mathbf{W}_{i}}$ and $\mathbf{\mathbf{o}_{j}}$ appear in the outer product matrix. Similarly, we define outer division, addition, and subtraction as
$(\mathbf{\mathbf{W}}_{i0}\oplus\mathbf{o}_{j0})_{ij}$ $=$ $\mathbf{W}_{0i}+\mathbf{o}_{0j}$ (4)
$(\mathbf{\mathbf{W}}_{i0}\ominus\mathbf{o}_{j0})_{ij}$ $=$ $\mathbf{W}_{0i}-\mathbf{o}_{0j}$ (5)
$(\mathbf{\mathbf{W}}_{i1}\oslash\mathbf{o}_{j1})_{ij}$ $=$ $\mathbf{W}_{1i}\div(\mathbf{o}_{1j}+\epsilon)$ (6)
where $\epsilon$ is a small number, and $\mathbf{W}_{i0}=[0;\mathbf{W}_{i}]$ and $\mathbf{o}_{j0}=[0;\mathbf{o}_{j}]$.
@@
The four matrices produced by MOAB are concatenated along the channel dimension to form a multimodal tensor $\mathbf{L}_{ij}\in\mathbb{R}^{4\times 257\times 257}$. We hypothesize that channel fusion will maintain the proximity of closer points and will use fewer parameters compared to a typical concatenation. By combining features across the channel dimension, we greatly decrease the dimension of the following fully connected layer from $(257x257)^{4}$ to $(257)^{2}$. Following the same parameters in [28], a 2D convolution layer is subsequently performed to leverage associated interactions, resulting in a singular condensed multimodal feature tensor $L_{ij}^{*}\in\mathbb{R}^{1\times 257\times 257}$. Last, we flatten $\mathbf{L}_{ij}^{*}$ and apply leaky ReLU followed by a linear predictor for brain tumor subtyping prediction. We hypothesize that MOAB simple yet effective operations adapt similar interactions conducted in [13] and rather excel fusion methods in [13] without introducing any approximation method.
@@ 2.2.4 Interpretability
Employing ABMIL on the post-fusion embedding, $\mathbf{p}_{ij}$ enables the visualization of attention scores, which provide an enriched perspective on the tumor’s phenotype and genotype. This, in turn, enhances ABMIL’s capacity to prioritize patches that are biologically relevant. Inspired by [13], we define our gated attention computation as follows:
@@
$\mathbf{W}_{i}$ $=f_{\rho}\left(\sum_{j=1}^{j}\mathbf{a}_{ij}\mathbf{h}_{ij}\right),$
Here, $\mathbf{W}_{p}$ and $\mathbf{W}_{g}$ are learnable parameters for the input and gating functions, respectively, with $\mathbf{b_{p}}$ and $\mathbf{b}_{g}$ as their corresponding biases. The symbol $\odot$ denotes element-wise multiplication, and $\sigma$ represents the sigmoid activation function. The gated embedding for patch $j$ in sample $i$ is given by $\mathbf{h}_{ij}$, while $\mathbf{a}_{ij}$ represents the attention weight for patch $j$, normalized across all patches. We leverage $\mathbf{a}_{ij}$ at inference to generate the heatmap shown in Fig. 5. To further enhance our late fusion stage, we weight $\mathbf{W}_{i}$ with $\mathbf{a}_{ij}$ using an MLP $f_{\rho}$ which consists of a linear transformation, an activation function (ReLU), and dropout. $f_{\rho}$ is applied to the pooled embedding $\mathbf{W}_{i}$ increases its representational power, making it a more refined input for the late fusion stage. This enriched embedding encapsulates a high-level representation of the WSI, effectively integrating morphological and molecular insights from the attention mechanism.
@@ 3.1 Performance Metrics
For the subtyping task, we assessed performance utilizing various metrics: F1-Macro, F1-Micro, Precision, and Recall/Sensitivity. F1-Macro is a significant metric for our quantitative analysis as it independently computes the F1 score for each class and subsequently averages them, assigning equal weight to each class irrespective of its size, thereby ensuring that the performance of minority classes is not overwhelmed by that of majority classes. For the survival prediction task, we chose performance metrics comparable to related work. Following an identical implementation to [13, 20, 22], For the survival prediction task, we define survival analysis as estimating the probability of an event occurring within a given survival time, while accounting for right-censored data. Censorship status is represented as $c\in\{0,1\}$, where $c=0$ denotes an observed event (e.g., death) and $c=1$ indicates the patient’s last known follow-up. Following previous work [13, 22, 20], we discretize the time-to-event into non-overlapping time intervals $(t_{i-1},t_{i}]$, based on the quartiles of survival times, and denote it as $y_{i}$. This formulation transforms the problem into a classification task with censorship information, where each patient is represented by $(\mathbf{L}_{\text{logits}},y_{i},c)$.
Next, we use the dual fusion embedding generated by MOAB, $\mathbf{L}_{\text{logits}}$, to predict the discretized bin corresponding to a time interval $t_{i}$. We define the discrete hazard function as:
@@
where $\sigma$ is the sigmoid activation function. Intuitively, $f_{\text{hazard}}(y_{i}\mid\mathbf{L}_{\text{logits}})$ represents the probability that the patient experiences the event (e.g., death) during the interval $(t_{i-1},t_{i}]$.
The discrete survival function is then defined as:
@@
which represents the probability that the patient survives up to the interval $(t_{i-1},t_{i}]$.
The Negative Log-Likelihood (NLL) survival loss is formally defined as:
($\mathcal{L}\left(\{\mathbf{L}_{\text{logits}},y,c\}_{i=1}^{N_{\text{total}}}\right)=-\sum_{i=1}^{N_{\text{total}}}\Big{[}c_{i}\log\big{(}f_{\text{surv}}(y_{i}\mid\mathbf{L}_{\text{logits}})\big{)}\\ +(1-c_{i})\log\big{(}f_{\text{surv}}(y_{i}\mid\mathbf{L}_{\text{logits}})-1[y_{i}]\big{)}\\ +(1-c_{i})\log\big{(}f_{\text{hazard}}(y_{i}\mid\mathbf{L}_{\text{logits}})\big{)}\Big{]}.$ (8)
where $(N_{\text{total}})$ is the total number of samples in the dataset, and $(k)$ corresponds to the total number of discretized labels. As noted in[13], this approach enhances survival analysis by accurately estimating survival probabilities, predicting event timestamps, and rigorously accounting for censorship status.
Note that in the ACMIL model proposed by Zhang et al. [23], the authors employed a composite loss function that integrates cross-entropy loss with a diversity loss designed to enhance the variability of the attention distribution across various regions of the WSI image. To modify ACMIL for survival analysis, we integrated the NLL loss with the diversity loss to emulate the learning methodology employed in the original study [23].
3.2 Domain Specific Prior Work and Ablation Studies
To evaluate our method on subtyping and survival prediction tasks, we replicate and adapt recent SOTA methods by incorporating MOAB as a replacement for the fusion technique initially used in these methods. For the survival prediction task, we conducted a comparative analysis using a consistent feature extractor across all modalities, including WSI and omic data, utilising the recent TCGA ID samples from [13]. We uniformly implemented training hyperparameters and loss functions across all models displayed in Tables 1 and 3. To this end, this section is divided into two parts. SOTA baseline models detailing unimodal and fusion models employed for each task, together with ablation studies included in Tables 1, 3, and 4.
@@ 3.2.1 State-of-the-Art Baseline Models
We evaluate our approach against SOTA techniques in the: 1) Unimodal and 2) Multimodal setting for both tasks, subtyping and survival prediction.
For the subtyping task:
@@
  • Unimodal techniques, we utilize MLP and SNN [32] as baseline models for the DNA methylation data. For WSI, we utilize Attention-based Multiple Instance Learning (ABMIL) [18], which implements gated weighted attention pooling to determine the importance of patches, and Transformer-based Multiple Instance Learning (TransMil) [31], which employs self-attention mechanisms from Transformers to evaluate correlations among patches in WSI. Through rigorous experimentation, we found that ABMIL outperforms the TransMIL baseline in both performance and computational efficiency, making it the preferred baseline for MOAD-Net, as also noted in [34].
@@
  • Multimodal techniques, we adapt Attention Challenging Multiple Instance Learning (ACMIL) [23] to work in a multimodal setting. ACMIL [23] is an approach proposed by Zhang et al. to address overfitting issues in MIL for single modality WSI classification. It employs multiple attention branches to identify varied discriminative patterns among instances, thereby distributing attention across the WSI and avoiding excessive concentration on a limited set of critical patches. Furthermore, TransMIL [31], with two late fusion variants, concatenation (Cat) and Kronecker product (KP), was conducted. We also utilize MCAT [10] and SurvPath [13], both of which perform multimodal tokenization to extract histology and biological pathway tokens. These tokens are then processed through a co-attention Transformer to model interactions between the different modalities. MCAT and SurvPath represent transformer-based early fusion approaches that integrate multimodal features at an initial stage. To maintain architectural consistency, we opted not to replace concatenation with the MOAB late fusion block in these models. The complex architectures of MCAT and SurvPath could be destabilized by MOAB, potentially introducing noise without improving performance. However, to test our dual fusion hypothesis, we replicated these models with a $\mathbf{p}_{ij}$ early fusion input representation, enabling each model to leverage two distinct fusion stages. By comparing the performance of SurvPath and MCAT with $\mathbf{p}_{ij}$ to their single-stage counterparts (SurvPath* and MCAT*), we provide empirical evidence that dual fusion outperforms single-stage fusion methods.
For the survival prediction task:
@@
  • Unimodal techniques, in addition to aforementioned unimodal models implemented for subtyping, we employed Spars-MLP [13], which tokenized transcriptomics into biological pathway tokens that encode specific cellular functions for the downstream analysis.
@@
  • Multimodal techniques, we integrated MOAD-Net across the same models used in subtyping tasks excluding MCAT, and including prototypical information bottlenecking and disentangling (PIBD) [20] method. PIBD [20] uses a prototypical information bottleneck to compress and align representations from different modalities, reducing complementary information redundancy while keeping the expressiveness of modality-specific information. It also introduces a disentanglement mechanism to separate modality-specific (WSI, omic) and shared information to integrate multimodal data. For PIBD experiments, we restricted MOAB to a late fusion setting, respecting PIBD’s initial modality-specific separation. Introducing MOAD-Net in an early fusion setting could potentially conflict with this initial modality independence, underscoring our strategic late fusion choice to maintain alignment with PIBD’s architectural principles.

### v2

@@ 2.2.1 Omic Encoder
To construct the omic encoder, we followed the established practice of using a Self-Normalizing Neural (SNN) [34] consisting of two fully connected layers, where each layer applies an Exponential Linear Unit (ELU) activation function followed by Alpha Dropout (0.25). The SNN compresses the 8K CpG sites into an encoded omic feature ${\mathbf{o}_{i}}\in\mathbb{R}^{d^{o}}$ ${d^{o}}=256$. Note that for fair comparison with other methods [22, 13, 20, 18], we used the same encoder tokenizing the genes into a group of pathways.
@@ 2.2.2 Whole Slide Image Encoder
For a WSI $\mathbf{X}$, we extract a collection of patches, represented as ${\mathbf{x}_{i0},\mathbf{x}_{i1},\ldots,\mathbf{x}_{ij}}$, where $i$ represents the patient/slide index and $j$ indicates the patch index that varies across slides. We extract non-overlapping patches from tissue areas at a 20× magnification (about 0.5 µm/pixel resolution). Subsequently, we utilized a SOTA image-only encoder [11] UNI to obtain the patch embeddings ${\mathbf{e}_{ij}}$. Using UNI ${f_{enc}(.)}$, we derive a set of low-dimensional patch embeddings for each patient, where ${\mathbf{e}_{ij}}={f_{enc}}(\mathbf{x}_{ij})\in\mathbb{R}^{d^{e}}$, ${d^{e}}=1024$, serving as input to our pipeline.
@@ 2.2.3 Fusion Stages
Early Fusion. Given a matrix of patch embeddings ${\mathbf{e}_{ij}}$ and the omic feature vector ${\mathbf{o}_{i}}$ for patient $i$, the encoded omic feature vector is cloned to match the number of patches in the WSI. This results in a tensor of shape $(N_{i},d_{o})$, where $N_{i}$ is the number of patches in WSI $i$, and $d_{o}$ is the dimension of the omic feature. The combined feature set for WSI $i$ is represented as:
@@
This operation is performed for all patches $j$ of the $i$th WSI. By incorporating omic features $\mathbf{o}_{i}$ into patch embeddings via early fusion, we enrich the representation with complementary molecular information. Leveraging an Attention-based Deep Multiple Instance Learning (ABMIL) approach[35], we capture patch-level discriminative features that synergistically combine molecular and morphological insights, enabling more precise identification of critical regions. The resulting embedding is projected to a slide-level representation $\mathbf{v}_{i}\in\mathbb{R}^{256}$, dimensionally aligned with the original omic feature $\mathbf{o}_{i}$ to facilitate subsequent late fusion.
@@
Late Fusion, motivated by [13] in modeling the interaction between omic to histology, histology to omic, and omic to omic, we mimic a similar behavior by employing our novel multimodal outer arithmetic fusion block (MOAB) [28] within a dual fusion approach, marking a new direction in combining MOAB with dual fusion to enhance modality integration. MOAB inputs ($\mathbf{v}_{i}$,$\mathbf{o}_{i}$) will be fused through four operations: outer product, outer division, outer subtraction, and outer addition. MOAB extracts various interactions while simultaneously preserving the $\mathbf{v}_{i}$ input feature by appending one to each input embedding when performing the outer product and division fusion and zero in the case of the outer subtraction and addition fusion. To simplify this, given the two embeddings $\mathbf{v}_{i}$, $\mathbf{o}_{i}$, $\mathbf{\mathbf{v}_{i}}\in\mathbb{R}^{N\times 1}$ and $\mathbf{\mathbf{o}_{i}}\in\mathbb{R}^{M\times 1}$. Initially, we append a 1 to each embedding, i.e., $\mathbf{v}_{i1}=[1;\mathbf{v}_{i}]$ and $\mathbf{o}_{j1}=[1;\mathbf{o}_{j}]$. Their outer product is defined as
$(\mathbf{\mathbf{v}}_{i1}\otimes\mathbf{o}_{j1})_{ij}=\mathbf{v}_{1i}*\mathbf{o}_{1j},$ (3)
@@
for $i\in[1...N+1]$ and $j\in[1...M+1]$ yielding a $(N+1)\times(M+1)$ matrix that intermingles every element of $\mathbf{v}_{i1}$ with every element of $\mathbf{o}_{j1}$. The appended 1 in both $\mathbf{v}_{i1}$ and $\mathbf{o}_{j1}$ ensures the original unimodal features $\mathbf{v}_{i}$ and $\mathbf{o}_{j}$ appear in the outer product matrix. Similarly, we define outer division, addition, and subtraction as
$(\mathbf{\mathbf{v}}_{i0}\oplus\mathbf{o}_{j0})_{ij}$ $=$ $\mathbf{v}_{0i}+\mathbf{o}_{0j}$ (4)
$(\mathbf{\mathbf{v}}_{i0}\ominus\mathbf{o}_{j0})_{ij}$ $=$ $\mathbf{v}_{0i}-\mathbf{o}_{0j}$ (5)
$(\mathbf{\mathbf{v}}_{i1}\oslash\mathbf{o}_{j1})_{ij}$ $=$ $\mathbf{v}_{1i}\div(\mathbf{o}_{1j}+\epsilon)$ (6)
where $\epsilon$ is a small number (set to $1\text{e}{-10}$), and $\mathbf{v}_{i0}=[0;\mathbf{v}_{i}]$ and $\mathbf{o}_{j0}=[0;\mathbf{o}_{j}]$.
@@
The four matrices produced by MOAB are concatenated along the channel dimension to form a multimodal tensor $\mathbf{L}\in\mathbb{R}^{4\times 257\times 257}$. We hypothesize that channel fusion will maintain the proximity of closer points and will use fewer parameters compared to a typical concatenation. By combining features across the channel dimension, we greatly decrease dimension by compressing the feature representation from $(257\times 257)^{4}$ to $(257)^{2}$. Following the same parameters in [28], a 2D convolution layer is subsequently performed to leverage associated interactions, resulting in a singular condensed multimodal feature tensor $\mathbf{L}^{*}\in\mathbb{R}^{1\times 257\times 257}$. Last, we flatten $\mathbf{L}^{*}$ and apply leaky ReLU followed by a linear predictor for brain tumor subtyping prediction. MOAB provides simple, yet effective operations to fuse multimodal data, similar to [13] but without relying on approximations.
@@ 2.2.4 Interpretability
Employing ABMIL [18] on the post-fusion embedding $\mathbf{p}_{ij}$ enables the visualization of attention scores, which provide an enriched perspective on the tumor’s visual and omics characterization. This, in turn, enhances ABMIL’s capacity to prioritize patches that are biologically relevant. We define our gated attention computation as follows:
@@
$\mathbf{v}_{i}$ $=f_{\rho}\left(\sum_{j=1}^{j}\mathbf{a}_{ij}\mathbf{h}_{ij}\right),$
Here, $\mathbf{W}_{p}$ and $\mathbf{W}_{g}$ are learnable parameters for the input and gating functions, respectively, with $\mathbf{b_{p}}$ and $\mathbf{b}_{g}$ as their corresponding biases. The symbol $\odot$ denotes element-wise multiplication, and $\sigma$ represents the sigmoid activation function. The gated embedding for patch $j$ in slide $i$ is given by $\mathbf{h}_{ij}$, while $\mathbf{a}_{ij}$ represents the attention weight for patch $j$, normalized across all patches. We leverage $\mathbf{a}_{ij}$ at inference to generate the heatmap shown in Fig. 3. To further enhance our late fusion stage, we weight $\mathbf{v}_{i}$ with $\mathbf{a}_{ij}$ using an MLP $f_{\rho}$ which consists of a linear transformation, an activation function (ReLU), and dropout. $f_{\rho}$ is applied to the pooled embedding $\mathbf{v}_{i}$ increases its representational power, making it a more refined input for the late fusion stage. This enriched embedding encapsulates a high-level representation of the WSI, effectively integrating morphological and molecular insights from the attention mechanism.
@@ 3.1 Performance Metrics
For the subtyping task, we assessed performance utilizing various metrics: F1-Macro, F1-Micro, Precision, and Recall/Sensitivity. F1-Macro is a significant metric for our quantitative analysis as it independently computes the F1 score for each class and subsequently averages them, assigning equal weight to each class irrespective of its size, thereby ensuring that the performance of minority classes is not overwhelmed by that of majority classes. For the survival prediction task, we follow the implementations of [13, 20, 22] where survival analysis is defined as an estimation of the probability of an event occurring within a given survival time, while accounting for right-censored data. Censorship status is represented as $c\in\{0,1\}$, where $c=0$ denotes an observed event (e.g., death) and $c=1$ indicates the patient’s last known follow-up. In line with previous work we discretize the time-to-event into non-overlapping time intervals $(t_{i-1},t_{i}]$, based on the quartiles of survival times denoted as $y_{i}$. This formulation transforms the problem into a classification task with censorship information, where each patient is represented by $(\mathbf{L}_{\text{logits}},y_{i},c)$.
Next, we use the dual fusion embedding generated by MOAB-FNet, $\mathbf{L}_{\text{logits}}$ to predict the discretized bin corresponding to a time interval $t_{i}$. We define the discrete hazard function as:
@@
where $\sigma$ is the sigmoid activation function. Intuitively, $f_{\text{hazard}}(y_{i}\mid\mathbf{L}_{\text{logits}})$ represents the probability that the patient experiences the event (e.g., death) during the interval $(t_{i-1},t_{i}]$. The discrete survival function is then defined as:
@@
which represents the probability that the patient survives up to the interval $(t_{i-1},t_{i}]$. The Negative Log-Likelihood (NLL) survival loss is formally defined as:
($\mathcal{L}\left(\{\mathbf{L}_{\text{logits}},y,c\}_{i=1}^{N_{\text{total}}}\right)=-\sum_{i=1}^{N_{\text{total}}}\Big{[}c_{i}\log\big{(}f_{\text{surv}}(y_{i}\mid\mathbf{L}_{\text{logits}})\big{)}\\ +(1-c_{i})\log\big{(}f_{\text{surv}}(y_{i}\mid\mathbf{L}_{\text{logits}})-1[y_{i}]\big{)}\\ +(1-c_{i})\log\big{(}f_{\text{hazard}}(y_{i}\mid\mathbf{L}_{\text{logits}})\big{)}\Big{]},$ (8)
where $(N_{\text{total}})$ is the total number of samples in the dataset, and $(k)$ corresponds to the total number of discretized labels.
3.2 Domain-Specific Prior Work and Ablation Studies
To evaluate our method on subtyping and survival prediction tasks, we replicate and adapt recent SOTA methods by incorporating MOAB as a replacement for the fusion technique initially used in these methods. For the survival prediction task, we conducted a comparative analysis using a consistent feature extractor across all modalities, including WSI and omic data, utilising the recent TCGA ID samples from [13]. We uniformly implemented training hyperparameters and loss functions across all models displayed in Tables 1 and 3. To this end, this section is divided into two parts. SOTA baseline models detailing unimodal and fusion models employed for each task, together with a description of the ablation studies conducted.
@@ 3.2.1 State-of-the-Art Baseline Models
We evaluate our approach against SOTA techniques in the unimodal and multimodal setting for both subtyping and survival prediction tasks.
@@
Unimodal Subtyping baselines. We utilize MLP and SNN [34] as baseline models for the DNA methylation data. For WSIs, we utilize Attention-based Multiple Instance Learning (ABMIL) [18], which implements gated weighted attention pooling to determine the importance of patches, as well as Transformer-based Multiple Instance Learning (TransMIL) [33] which employs the Nyström attention mechanism to evaluate correlations among WSI patches. Through rigorous experimentation, we found that ABMIL outperforms the TransMIL baseline in both performance and computational efficiency, also noted in [36], making it the preferred baseline for MOAD-FNet.
@@
Multimodal Subtyping baselines. We adapt Attention Challenging Multiple Instance Learning (ACMIL) [23] to work in a multimodal setting. ACMIL is an approach designed to address overfitting in single-modality WSI classification, by using multiple attention branches and a composite loss (cross-entropy + diversity loss) to distribute attention across the WSI. We also used TransMIL with two late fusion variants, concatenation (Cat) and Kronecker product (KP) [33]. Furthermore, we compare against MCAT [10] and SurvPath [13], both of which perform multimodal tokenization to extract histology and biological pathway tokens. MCAT and SurvPath employ a transformer-based early fusion approach, the concatenate the resulting vectors. To maintain their architectural consistency, we opted not to replace concatenation with the MOAB late fusion block, as their complex architectures could be destabilized by it. However, to test our dual fusion hypothesis we input our early fusion input representation $\mathbf{p}_{ij}$ into MCAT and SurvPath, thus enabling these models to leverage omic-WSIs embeddings at both the early and late fusion stages. By comparing the performance of SurvPath and MCAT with $\mathbf{p}_{ij}$ to the original implementation (SurvPath* and MCAT*), we provide evidence that dual fusion outperforms single-stage fusion methods.
@@
Unimodal Survival baselines. In addition to the unimodal subtyping baselines, we employ Sparse-MLP [13], which tokenizes transcriptomics into biological pathway tokens encoding specific cellular functions for the downstream analysis.
@@
Multimodal Survival baselines. We integrated MOAD-FNet across the same models used for subtyping tasks, adding the prototypical information bottlenecking and disentangling (PIBD) method [20]. It introduces a disentanglement mechanism to separate modality-specific versus shared information. For PIBD, we restricted MOAB to a late fusion setting, respecting PIBD’s initial modality-specific separation.